OpenStudy (anonymous):

Solve each absolute value inequality and graph the solution set. 3|4q-2| less than or equal to 18

OpenStudy (anonymous):

@mathmath333 @kisstherains @blackops2luvr @DarkBlueChocobo @data_lg2 @sammixboo @ganeshie8

OpenStudy (anonymous):

Math isn't my best subject..sorry.

OpenStudy (anonymous):

@blackops2luvr it says the page can not be found

OpenStudy (blackops2luvr):

ugh look up ur question and find the \(\normalsize\sf\color{black}{Link}\)

OpenStudy (anonymous):

like on Google?

OpenStudy (blackops2luvr):

\(\normalsize\sf\color{purple}{yes}\)

sammixboo (sammixboo):

I forgot how to do this o _ o I am so sorry D:

OpenStudy (anonymous):

:( okay thanks anyways @sammixboo

OpenStudy (anonymous):

@blackops2luvr or @mathmath333 could you help? :/

OpenStudy (anonymous):

How do I graph it ? and q would be less than or equal to -2 right?

OpenStudy (mathmath333):

i think u have to find q through this since absolute value is asked \(\large\tt \begin{align} \color{black}{-18\leq 3(4q-2) \leq18}\end{align}\)

OpenStudy (anonymous):

How would I do that ?

OpenStudy (anonymous):

so does the first part of what you said not included in how to find the absolute value?

OpenStudy (mathmath333):

yes it would be better to do with i did knoe

OpenStudy (anonymous):

* absolute value inequality

OpenStudy (mathmath333):

*now

OpenStudy (anonymous):

SO how do I find the absolute value inequality of the equation then?

OpenStudy (mathmath333):

first solve for q

OpenStudy (anonymous):

I got -2

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{-18\leq 3(4q-2) \leq18\\ -6\leq (4q-2) \leq6\\ -4\leq 4q \leq8\\ -1\leq q \leq2\\}\end{align}\)

OpenStudy (anonymous):

What is that for?

OpenStudy (anonymous):

Is that the inequality?

OpenStudy (mathmath333):

wait

OpenStudy (anonymous):

okay

OpenStudy (mathmath333):



OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ -1\leq q \leq2\\}\end{align}\) we can graph this on number line

OpenStudy (anonymous):

okay. so the answer is what you just wrote and then we just graph it ?

OpenStudy (mathmath333):

solution set is this \(\Large\tt \begin{align} \color{black}{[-1,2]}\end{align}\)

OpenStudy (anonymous):

okay and the absoulte value inequality is that -2?

OpenStudy (mathmath333):

wait the graph of that will be this http://symbolab.com/solver/absolute-inequalities-calculator/3%7C4q-2%7C%20%5Cle18/?origin=button

OpenStudy (mathmath333):

a triangle is formed